OpenStudy (anonymous):

cos(a/2) = + or - sqrt(1 + cos a / 2 ) for all values of a. True or false?

OpenStudy (anonymous):

write it this way: cos(a/2)^2=(1+cos a)/2 then: 2 cos^2(a/2)=(1+cos a) now: cos a = cos^2(a/2)-sin^2(a/2) and 1 =sin^2(a/2) +cos^2(a/2) so: 1+cos a = sin^2(a/2) +cos^2(a/2) + cos^2(a/2)-sin^2(a/2) =2cos^2(a/2) so it is true

OpenStudy (anonymous):

@Spectrum